wym no idea >:( let someone else try it thenDetermine whether a conclusion can be drawn about the existence of uniqueness of a solution of the differential equation (8 -
Answer #1

SOLUTION = (8-4) 9+ 3t y_ 4y = sin() y (o)= 8. YO)= 8 DIVIDED BY (8-4) Sinio 8- t 94(34)99- 1)4-6644)y=- 9+(134) 94ön)y= s